Given the following National Income Model

Y = C + I0 + G0 + X0 – Z
Where C = C0 + bYd
Yd = Y – T

T = T0 + Ty and Z = Z0 + ZYd
In this equation X and Z are export and import
• Identify endogenous and exogenous variables and parameters.
• Find out equilibrium level of Y and C
(full solution)
1
Expert's answer
2020-06-26T09:33:02-0400

1) Identify endogenous and exogenous variables and parameters


The exogenous variables are "I_0, \;G_0, \; X_0, \; T_0, \; Z_0"


The endogenous variables are "Y, \; C, \; T, \; Z"



2) Find out equilibrium level of Y and C


At equilibrium,

"Y = C + I + G + (X - Z)"

Therefore:



"Y = C_0 + b(Y - T_0 - tY) + I_0 + G_0 + X_0 - z(Y - T_0 - tY)\\[0.3cm]
Y(1 - b + tb + z - tz) = C_0 - bT_0 + I_0 + G_0 + X_0 + zT_0\\[0.3cm]
\color{red}{Y^* = \dfrac{C_0 - bT_0 + I_0 + G_0 + X_0 + zT_0}{1 - b + tb + z - tz}}"

This gives an equilibrium consumption of:


"C = C_0 + bY_d\\[0.3cm]
C = C_0 + b(Y - T_0 - tY)\\[0.3cm]
C^* = C_0 + b(1 - t)Y^* - bT_0\\[0.3cm]

C^* = C_0 + \dfrac{b(1 - t)(C_0 - bT_0 + I_0 + G_0 + X_0 + zT_0)}{1 - b + tb + z - tz} - bT_0\\[0.3cm]

\color{red}{C^* = \dfrac{(C_0 - bT_0)(1 - b + tb + z - tz) + b(1 - t)(C_0 - bT_0 + I_0 + G_0 + X_0 + zT_0)}{1 - b + tb + z - tz}}"